# Query, Don’t Train: Privacy-Preserving Tabular Prediction from EHR Data via SQL Queries

### v3

@@ 1 Introduction
Despite these constraints, public datasets such as MIMIC-III have enabled research in EHR-driven prediction under carefully controlled conditions, supporting tasks such as mortality forecasting, hospital readmission risk, and treatment efficacy modeling [11, 19]. Traditional supervised models—especially tree-based methods like XGBoost—continue to dominate tabular prediction tasks due to their robustness to heterogeneous features, irregular target functions, and missing data [8, 25, 18]. Transformer-based in-context learners, such as TabPFN, offer classification via training-set conditioning, though they still require access to raw examples at inference time [10, 6, 21, 2].
LLMs have recently demonstrated strong performance both in structured reasoning tasks, including text-to-SQL translation [7], and tabular prediction tasks [9, 13]. Recent advancements have introduced privacy-preserving techniques and agent-based frameworks to address these challenges [17]. Deep learning models can be trained across decentralized datasets using federated learning, enabling collaborative prediction without sharing sensitive data [1, 4, 16, 24]. Agent-based frameworks allow models to autonomously perform multi-step reasoning over structured data, facilitating complex clinical decision-making [17].
@@ 2.2 Datasets
We focus on predicting 30-day hospital readmissions for patients with Type 2 Diabetes in US hospitals [diabetes_130-us_hospitals_for_years_1999-2008_296] [^2] https://www.kaggle.com/c/1056lab-diabetes-readmission-prediction/data. The dataset consists of patient records $x_{i}$, which include demographics, laboratory results, procedures, and prior admissions, with binary outcome labels $y_{i}\in\{0,1\}$. We conducted K-fold cross-validation with 5 iterations, utilizing a subset of 2,000 randomly sampled patients in each fold. With approximately 12% of patients readmitted within 30 days in our evaluation subset, this approach effectively addresses the imbalanced nature of the readmission task.
@@ 2.4 Classification Results
XGBoost $0.65\pm\scriptstyle{0.04}$ $0.68\pm\scriptstyle{0.06}$ $0.66\pm\scriptstyle{0.05}$

### v4

@@ 1 Introduction
Despite these constraints, public datasets such as MIMIC-III have enabled research in EHR-driven prediction under carefully controlled conditions, supporting tasks such as mortality forecasting, hospital readmission risk, and treatment efficacy modeling [12, 20, 28]. Traditional supervised models—especially tree-based methods like XGBoost—continue to dominate tabular prediction tasks due to their robustness to heterogeneous features, irregular target functions, and missing data [9, 27, 19]. Transformer-based in-context learners, such as TabPFN, offer classification via training-set conditioning, though they still require access to raw examples at inference time [11, 7, 22, 2].
LLMs have recently demonstrated strong performance both in structured reasoning tasks, including text-to-SQL translation [8, 23], and tabular prediction tasks [10, 14]. Recent advancements have introduced privacy-preserving techniques and agent-based frameworks to address these challenges [18]. Deep learning models can be trained across decentralized datasets using federated learning, enabling collaborative prediction without sharing sensitive data [1, 4, 17, 26]. Agent-based frameworks allow models to autonomously perform multi-step reasoning over structured data, facilitating complex clinical decision-making [18].
@@ 2.2 Datasets
We focus on predicting 30-day hospital readmissions for patients with Type 2 Diabetes in US hospitals [5] [^2] https://www.kaggle.com/c/1056lab-diabetes-readmission-prediction/data. The dataset consists of patient records $x_{i}$, which include demographics, laboratory results, procedures, and prior admissions, with binary outcome labels $y_{i}\in\{0,1\}$. We conducted K-fold cross-validation with 5 iterations, utilizing a subset of 2,000 randomly sampled patients in each fold. With approximately 12% of patients readmitted within 30 days in our evaluation subset, this approach effectively addresses the imbalanced nature of the readmission task.
@@ 2.4 Classification Results
XGBoost $0.65\pm\scriptstyle{0.04}$ $8\pm\scriptstyle{0.06} 0.66±0$$6\pm\scriptstyle{0.05}$